given f(t)=2t+3, does the equation have no solution, one solution, or an infinite number of solutions?
f(t) + 1= f(t+1)

Answer by TimothyLamb(4379)

given:
f(t) = 2t + 3
---
f(t) + 1 = f(t+1)
2t + 3 + 1 = 2(t + 1) + 3
2t + 4 = 2t + 2 + 3
2t + 4 = 2t + 5
4 = 5
---
answer:
no solution
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---